# Multiclass Local Calibration With the Jensen-Shannon Distance

### v1

@@ B.1.1 Proof of Theorem 2
Choosing $\tau_{b}=\sqrt{\tfrac{\log(2Cm_{B}/\delta)}{2|B_{b}|}}$ and applying the union bound over the $m_{B}\cdot C$ bin–class pairs yields: with probability at least $1-\delta$,
$\forall b,c:\qquad\big|\mathrm{freq}_{b,c}-\hat{\theta}_{b,c}\big|\leq\sqrt{\frac{\log(2Cm_{B}/\delta)}{2|B_{b}|}}.$ (16)
@@
$\mathcal{E}(D;\varphi;\beta)$ $\leq L_{\varphi}\sum_{b=1}^{m_{B}}w_{b}\sum_{c=1}^{C}\pi_{c}\Big(\sqrt{\frac{\log(2Cm_{B}/\delta)}{2|B_{b}|}}+\varepsilon\Big)$
$=L_{\varphi}\Big(\varepsilon\sum_{c=1}^{C}\pi_{c}\sum_{b=1}^{m_{B}}w_{b}\;+\;\sum_{b=1}^{B}w_{b}\sqrt{\frac{\log(2Cm_{B}/\delta)}{2|B_{b}|}}\sum_{c}\pi_{c}\Big).$
Using $\sum_{b}w_{b}=1$ and $\sum_{c}\pi_{c}=1$, this simplifies to
$\boxed{\Pr\Bigg(\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\Big[\varepsilon+\sum_{b=1}^{m_{B}}w_{b}\sqrt{\frac{\log(2Cm_{B}/\delta)}{2|B_{b}|}}\Big]\Bigg)\geq 1-\delta.}$ (17)
Therefore achieving the form of theorem 2 when $\Psi(b,\mathcal{Y})=B_{b}$.
Confidence Based Binning
$\mathcal{E}(D;\varphi;\beta)$ $\leq L_{\varphi}\sum_{b=1}^{B}w_{b}\sum_{c=1}^{C}\pi_{c}\Big(\sqrt{\frac{\log(2Cm_{B}/\delta)}{2|B_{b,c}|}}+\varepsilon\Big)\leq L_{\varphi}\Bigg[\varepsilon+\sum_{b=1}^{B}w_{b}\sqrt{\frac{\log(2Cm_{B}/\delta)}{2\min_{c}|B_{b,c}|}}\Bigg]$ (18)
and by setting $\Psi(b,\mathcal{Y})=B_{b,c^{*}}$ and $c^{*}=\arg\min_{c\in\mathcal{Y}}|B_{b,c}|$ we obtain a bound of the same form of theorem 2. ∎
Proof of Corollary 1.
@@
$\big|\mathrm{conf}_{b,c}-\hat{\theta}_{b,c}\big|\leq\frac{1}{|B_{b}|}\sum_{i\in I_{b}}|\mathbf{\hat{p}}_{i,c}-\hat{\theta}_{c}(\mathbf{y}_{i}\mid\mathbf{x}_{i})|\leq\varepsilon.$
@@
$\Pr\Bigg(\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\Big[C\cdot L\cdot\rho+\sum_{b=1}^{m_{B}}w_{b}\sqrt{\frac{\log(2Cm_{B}/\delta)}{2|\Psi(b,\mathcal{Y})|}}\Big]\Bigg)\geq 1-\delta.$
@@
$\boxed{\limsup_{\rho\to 0}\Pr\Bigg(\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\sum_{b=1}^{m_{B}}w_{b}\sqrt{\frac{\log(2Cm_{B}/\delta)}{2|\Psi(b,\mathcal{Y})|}}\Bigg)\geq 1-\delta}.$ (7)
B.1.2 Proof of Theorem 3
@@
Let $f:\mathcal{X}\to\Delta^{C}$ be a probabilistic classifier, where $f$ is composed of a feature extractor $\phi:\mathcal{X}\to\mathcal{F}$ and a final classification layer $g:\mathcal{F}\to\Delta^{C}$. Assume $f$ to be locally calibrated up to error $\varepsilon$ and let $k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})$ kernel functions to obtain the kernel-weighted mean of both the empirical frequencies and the predicted probabilities for given anchor point $\mathbf{x}_{i}$:
@@
$\text{LCE}=\frac{1}{C}\sum_{b=1}^{B}\frac{|B_{b}|}{n}\frac{1}{|B_{b}|}\sum_{i\in b}\left\|\frac{\sum_{j\in b}\bigl(\mathbf{\hat{p}}_{j}-\mathbf{y}_{j}\bigr)\,k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})}{\sum_{j\in b}k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})}\right\|_{1}=\frac{1}{C}\sum_{b=1}^{B}\frac{|B_{b}|}{n}\frac{1}{|B_{b}|}\sum_{i\in b}\|\hat{\theta}(\mathbf{\hat{p}}_{i}\mid\mathbf{x}_{i})-\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})\|_{1}$
We can rewrite as follows:
$\frac{1}{|B_{b}|}\sum_{i\in b}\|\hat{\theta}(\mathbf{\hat{p}}_{i}\mid\mathbf{x}_{i})-\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})\|_{1}=\frac{1}{|B_{b}|}\sum_{i\in b}\|\hat{\theta}(\mathbf{\hat{p}}_{i}\mid\mathbf{x}_{i})-\mathbf{\hat{p}}_{i}-\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})+\mathbf{\hat{p}}_{i}\|_{1}\leq$
$\frac{1}{|B_{b}|}\sum_{i\in b}\big\|\hat{\theta}(\mathbf{\hat{p}}_{i}\mid\mathbf{x}_{i})-\mathbf{\hat{p}}_{i}\big\|_{1}+\frac{1}{|B_{b}|}\sum_{i\in b}\|\mathbf{\hat{p}}_{i}-\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})\|_{1}$
According to same local calibration assumption of Theorem 2, the last component in bounded as follows:
$\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\|\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})-\mathbf{\hat{p}}_{i}\|_{1}\leq\varepsilon.$
@@
$\|\hat{\theta}(\mathbf{\hat{p}}_{i}\mid\mathbf{x}_{i})-\mathbf{\hat{p}}_{i}\big\|_{1}=\Big\|\sum_{j\in I_{b}}w_{ij}\mathbf{\hat{p}}_{j}-\mathbf{\hat{p}}_{i}\Big\|_{1}=\Big\|\sum_{j\in I_{b}}w_{ij}(\mathbf{\hat{p}}_{j}-\mathbf{\hat{p}}_{i})+\sum_{j\in I_{b}}w_{ij}\mathbf{\hat{p}}_{i}-\mathbf{\hat{p}}_{i}\Big\|_{1}=$
$\Big\|\sum_{j\in I_{b}}w_{ij}(\mathbf{\hat{p}}_{j}-\mathbf{\hat{p}}_{i})+(\underbrace{\sum_{j\in I_{b}}w_{ij}}_{=1}-1)\mathbf{\hat{p}}_{i}\Big\|_{1}=\Big\|\sum_{j\in I_{b}}w_{ij}(\mathbf{\hat{p}}_{j}-\mathbf{\hat{p}}_{i})\Big\|_{1}$
@@
$\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\Big\|\sum_{j\in I_{b}}w_{ij}(\mathbf{\hat{p}}_{j}-\mathbf{\hat{p}}_{i})\Big\|_{1}\leq\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\underbrace{\sum_{j\in I_{b}}w_{i,j}\|\mathbf{\hat{p}}_{j}-\mathbf{\hat{p}}_{i}\|_{1}}_{Z_{i}}$
Again we follow the conditioning assumptions of Appendix A.2. Please note that each coordinate-wise distance satisfies $\|\mathbf{\hat{p}}_{j}-\mathbf{\hat{p}}_{i}\|_{1}\in[0,2]$, hence $Z_{i}\in[0,2]$. We can now apply the weighted version of Hoeffding’s inequality to each centered quantity $Z_{i}-\mathbb{E}[Z_{i}]$, conditioning on the kernel weights and features to obtain zero-mean summands. The union bound then yields a simultaneous statement over anchors. We proceed fixing a bin $b$ with index set $I_{b}$ of size $|B_{b}|$. For each anchor $i\in I_{b}$ define the effective sample size associated with the weights:
@@
$\Pr\Bigg(Z_{i}-\mathbb{E}[Z_{i}]\geq\tau\Bigg)\leq\text{exp}\Bigg(-\frac{n_{i}^{\text{eff}}\tau^{2}}{2}\Bigg)$
And, by setting $\delta^{\prime}=\delta/n$ we have:
$\Pr\Bigg(Z_{i}-\mathbb{E}[Z_{i}]\leq\sqrt{\frac{2\log\!\bigl(1/\delta^{\prime}\bigr)}{n_{i}^{\mathrm{eff}}}}\Bigg)\geq 1-\delta^{\prime}$
@@
$Z_{i}-\mathbb{E}[Z_{i}]\leq\sqrt{\frac{2\log\!\bigl(n/\delta\bigr)}{n_{i}^{\mathrm{eff}}}}$
@@
$\Pr\Bigg(\frac{1}{|B_{b}|}\sum_{i\in I_{b}}Z_{i}\leq\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\mathbb{E}[Z_{i}]+\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\sqrt{\frac{2\log\!\bigl(n/\delta\bigr)}{n_{i}^{\mathrm{eff}}}}\Bigg)\geq 1-\delta$
We conclude the proof by providing a bound for the expectation in the context of a neural network classifier $\phi(\cdot)$. More precisely, by the Lipschitz continuity of the softmax [4],
$\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\sum_{j\in I_{b}}w_{i,j}\|\mathbf{\hat{p}}_{j}-\mathbf{\hat{p}}_{i}\|_{1}\leq\frac{L}{|B_{b}|}\sum_{i\in I_{b}}\sum_{j\in I_{b}}w_{i,j}\|\phi(\mathbf{x}_{j})-\phi(\mathbf{x}_{i})\|_{1}.$

### v2

@@ B.2 Proof of Theorem 2
Choosing $\tau_{b}=\sqrt{\tfrac{\log(\frac{2Cm_{B}}{\delta})}{2|B_{b}|}}$ and applying the union bound over the $m_{B}\cdot C$ bin–class pairs yields: with probability at least $1-\delta$,
$\forall\;b,c:\qquad\big|\mathrm{freq}_{b,c}-\hat{\theta}_{b,c}\big|\leq\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2|B_{b}|}}.$ (18)
@@
$\boxed{\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\sum_{b=1}^{m_{B}}\sum_{c=1}^{C}w_{b}\cdot\pi_{c}\Bigg(\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2|B_{b}|}}+\varepsilon\Bigg)}$ (19)
Therefore achieving the form of Theorem 2 when $\Psi(b,c)=B_{b}$ and $w_{b,c}=w_{b}$.
Class-wise metrics.
In the context of class-wise metrics, the only difference is the introduction of a dependence between bins’ cardinalities and the classes: $|B_{b}|\to|B_{b,c}|$. Then, a bound of same style applies and with probability at least $1-\delta$:
$\mathcal{E}(D;\varphi;\beta)$ $\leq L_{\varphi}\sum_{b=1}^{m_{B}}\sum_{c=1}^{C}w_{b,c}\cdot\pi_{c}\Bigg(\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2|B_{b,c}|}}+\varepsilon\Bigg)$
and by setting $\Psi(b,c)=B_{b,c}$ we obtain a bound of the same form of Theorem 2. ∎
Discussion on finite-sample kernel estimates.
The proof above assumes that the reference kernel estimates are computed on an arbitrarily large disjoint set, so that they may be treated as deterministic. When the Nadaraya–Watson estimator is instead computed on a finite sample, additional stochastic error arises.
Fix a bin $b$, class $c$, and anchor point $\mathbf{x}_{i}$. The kernel estimator can be written as
$\hat{\theta}_{i,c}=\sum_{j\in I_{b}}w_{ij}\,\mathbf{1}\{y_{j}=c\},\qquad w_{ij}=\frac{k_{\gamma_{n}}(\mathbf{x}_{i},\mathbf{x}_{j})}{\sum_{\ell\in I_{b}}k_{\gamma_{n}}(\mathbf{x}_{i},\mathbf{x}_{\ell})}.$ (20)
Conditioning on the inputs $\{\mathbf{x}_{j}\}_{j\in I_{b}}$, the weights $w_{ij}$ are deterministic and the only randomness stems from the labels $\{y_{j}\}_{j\in I_{b}}$. Since the label indicators $\mathbf{1}\{y_{j}=c\}$ are independent and bounded in $[0,1]$, a weighted Hoeffding inequality applies. In particular,
$\Pr\!\left(\big|\hat{\theta}_{i,c}-\mathbb{E}[\hat{\theta}_{i,c}\mid X]\big|\geq\tau\;\middle|\;X\right)\leq 2\exp\!\left(-2\tau^{2}\,n_{i}^{\mathrm{eff}}\right),$ (21)
where the effective sample size is
$n_{i}^{\mathrm{eff}}:=\frac{1}{\sum_{j\in I_{b}}w_{ij}^{2}}.$ (22)
Applying a union bound over all $m_{B}C$ bin–class pairs yields that, with probability at least $1-\delta$,
$\forall b,c:\quad\big|\hat{\theta}_{b,c}-\mathbb{E}[\hat{\theta}_{b,c}\mid X]\big|\leq\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2\,n_{b}^{\mathrm{eff}}}},$ (23)
where $n_{b}^{\mathrm{eff}}$ denotes a lower bound on the effective sample size within bin $b$.
Now decompose
$\big|\mathrm{freq}_{b,c}-\hat{\theta}_{b,c}\big|=\big|\mathrm{freq}_{b,c}-\mathbb{E}[\hat{\theta}_{b,c}\mid X]-\hat{\theta}_{b,c}+\mathbb{E}[\hat{\theta}_{b,c}\mid X]\big|$ (24)
and apply the triangle inequality:
$\big|\mathrm{freq}_{b,c}-\hat{\theta}_{b,c}\big|\leq\big|\mathrm{freq}_{b,c}-\mathbb{E}[\hat{\theta}_{b,c}\mid X]\big|+\big|\hat{\theta}_{b,c}-\mathbb{E}[\hat{\theta}_{b,c}\mid X]\big|.$ (25)
The first term is the standard empirical fluctuation term and satisfies
$\big|\mathrm{freq}_{b,c}-\mathbb{E}[\hat{\theta}_{b,c}\mid X]\big|\leq\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2\,|B_{b}|}}.$ (26)
The second term satisfies
$\big|\hat{\theta}_{b,c}-\mathbb{E}[\hat{\theta}_{b,c}\mid X]\big|\leq\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2\,n_{b}^{\mathrm{eff}}}}.$ (27)
Since $n_{b}^{\mathrm{eff}}\leq|B_{b}|$, the two terms are of the same order and we may upper bound their sum by
$\big|\mathrm{freq}_{b,c}-\hat{\theta}_{b,c}\big|\leq 2\;\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2\,n_{b}^{\mathrm{eff}}}}.$ (28)
Moreover, under standard smoothness assumptions on the conditional class probabilities (Appendix A.3), the kernel-smoothed conditional probability differs from the true conditional probability by a bias of order $\mathcal{O}(\gamma_{n}^{2})$ (Wasserman, 2006). Combining stochastic concentration and smoothing bias yields
$\big|\mathrm{freq}_{b,c}-\hat{\theta}_{b,c}\big|\leq 2\;\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2\,n_{b}^{\mathrm{eff}}}}+\mathcal{O}(\gamma_{n}^{2}).$ (29)
Therefore the bound in Eq. 19 becomes:
$\Pr\Bigg(\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\sum_{b=1}^{m_{B}}\sum_{c=1}^{C}w_{b}\cdot\pi_{c}\Big(\varepsilon+2\;\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2\,n_{b}^{\mathrm{eff}}}}+\mathcal{O}(\gamma_{n}^{2})\Big)\Bigg)\geq 1-\delta.$ (30)
The additional $\mathcal{O}(\gamma_{n}^{2})$ term vanishes as $n\to\infty$ under the kernel consistency assumptions.
B.3 Proof of Corollary 1.
Statement. If a classifier $f$ satisfies $\rho$-perfect local calibration, the error $\mathcal{E}(D;\varphi;\beta)$ becomes purely stochastic fluctuation. For any $\delta\in[0,1]$, with probability at least $1-\delta$ it holds that:
$\limsup_{\rho\to 0}\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\sum_{b=1}^{m_{B}}\sum_{c=1}^{C}w_{b,c}\,\cdot\,\pi_{c}\,\,\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2|\Psi(b,c)|}}$ (31)
@@
$\big|\mathrm{conf}_{b,c}-\hat{\theta}_{b,c}\big|\leq\frac{1}{|B_{b,c}|}\sum_{i\in I_{b,c}}|\mathbf{\hat{p}}_{i,c}-\hat{\theta}_{c}(\mathbf{y}_{i}\mid\mathbf{x}_{i})|\leq\varepsilon.$
@@
$\Pr\Bigg(\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\sum_{b=1}^{m_{B}}\sum_{c=1}^{C}w_{b,c}\cdot\pi_{c}\Bigg(\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2|\Psi(b,c)|}}+C\cdot L\cdot\rho\Bigg)\Bigg)\geq 1-\delta.$
@@
$\boxed{\limsup_{\rho\to 0}\Pr\Bigg(\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\sum_{b=1}^{m_{B}}\sum_{c=1}^{C}w_{b,c}\cdot\pi_{c}\,\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2|\Psi(b,c)|}}\Bigg)\geq 1-\delta}.$ (7)
B.4 Proof of Theorem 3
Statement. Let $f$ be composed of a feature extractor $\phi$ and classification layer $g$. Let $f$ be locally calibrated with error $\varepsilon$ relative to kernel estimates $\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})$ computed on an independent arbitrarily large auxiliary dataset. Let the LCE be computed on $\phi(\mathbf{x}_{i})$. Then, conditioning on the evaluation inputs $\{\mathbf{x}_{i}\}_{i=1}^{n}$ and on the corresponding consistent kernel estimates, for any $\delta\in[0,1]$, with at least probability $1-\delta$:
$\mathrm{LCE}\leq\;\varepsilon\;+\;\frac{2}{n}\sum_{b=1}^{m_{B}}\sum_{i\in I_{b}}\sqrt{\frac{\log\!\bigl(\frac{2nC}{\delta}\bigr)}{2n_{i}^{\mathrm{eff}}}}\>+\>\frac{L}{n}\sum_{b=1}^{m_{B}}\mathbb{E}\!\Big[\sum_{i\in I_{b}}\sum_{j\in I_{b}}w_{i,j}\|\phi(\mathbf{x}_{j})-\phi(\mathbf{x}_{i})\|_{1}\Big]s$
@@ B.4 Proof of Theorem 3
Let $k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})$ kernel functions to obtain the kernel-weighted mean of both the empirical frequencies and the predicted probabilities for given anchor point $\mathbf{x}_{i}$:
@@
$\text{LCE}=\frac{1}{C}\sum_{b=1}^{m_{B}}\frac{|B_{b}|}{n}\frac{1}{|B_{b}|}\sum_{i\in b}\left\|\frac{\sum_{j\in b}\bigl(\mathbf{\hat{p}}_{j}-\mathbf{y}_{j}\bigr)\,k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})}{\sum_{j\in b}k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})}\right\|_{1}=\frac{1}{C}\sum_{b=1}^{m_{B}}\frac{|B_{b}|}{n}\frac{1}{|B_{b}|}\sum_{i\in b}\|\hat{\theta}(\mathbf{\hat{p}}_{i}\mid\mathbf{x}_{i})-\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})\|_{1}$
In the following, we bound the deviation coordinate-wise and obtain a uniform statement over all classes $c\in[C]$. Fix a bin $b$ and consider one coordinate $c$. By triangle inequality,
$\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\big|\hat{\theta}(\hat{p}_{i,c}\mid\mathbf{x}_{i})-\hat{\theta}(y_{i,c}\mid\mathbf{x}_{i})\big|$ $\leq\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\big|\hat{\theta}(\hat{p}_{i,c}\mid\mathbf{x}_{i})-\hat{p}_{i,c}\big|+\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\big|\hat{p}_{i,c}-\hat{\theta}(y_{i,c}\mid\mathbf{x}_{i})\big|$
By local calibration with respect to the (infinite-sample) kernel target $\theta(\cdot\mid\mathbf{x}_{i})$, $\|\hat{\mathbf{p}}_{i}-\theta(\mathbf{y}\mid\mathbf{x}_{i})\|_{1}\leq\varepsilon$, hence for every $c$, $|\hat{p}_{i,c}-\theta(y_{i,c}\mid\mathbf{x}_{i})|\leq\varepsilon$. Therefore, for each $i\in I_{b}$,
$\big|\hat{\theta}(y_{i,c}\mid\mathbf{x}_{i})-\hat{p}_{i,c}\big|\leq\big|\hat{\theta}(y_{i,c}\mid\mathbf{x}_{i})-\theta(y_{i,c}\mid\mathbf{x}_{i})\big|+\varepsilon$ (32)
Moreover, conditioning on the inputs, $\hat{\theta}(y_{i,c}\mid\mathbf{x}_{i})$ is a weighted average of bounded independent label indicators. Thus, we apply Hoeffding’s inequality and for each fixed $(i,c)$ and with probability at least $1-\delta^{\prime}$,
$\big|\hat{\theta}(y_{i,c}\mid\mathbf{x}_{i})-\theta(y_{i,c}\mid\mathbf{x}_{i})\big|\leq\sqrt{\frac{\log(\frac{2}{\delta^{\prime}})}{2\,n_{i}^{\mathrm{eff}}}},\qquad n_{i}^{\mathrm{eff}}:=\Big(\sum_{j\in I_{b}}w_{ij}^{2}\Big)^{-1}$ (33)
Setting $\delta^{\prime}=\delta/(nC)$ and applying a union bound over all anchors $i$ and classes $c$ implies that with probability at least $1-\delta$,
$\forall i,c:\quad\big|\hat{\theta}(y_{i,c}\mid\mathbf{x}_{i})-\theta(y_{i,c}\mid\mathbf{x}_{i})\big|\leq\sqrt{\frac{\log(\frac{2nC}{\delta})}{2\,n_{i}^{\mathrm{eff}}}}$ (34)
Averaging over $i\in I_{b}$ yields
$\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\big|\hat{\theta}(y_{i,c}\mid\mathbf{x}_{i})-\hat{p}_{i,c}\big|\leq\varepsilon+\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\sqrt{\frac{\log(\frac{2nC}{\delta})}{2n_{i}^{\mathrm{eff}}}}$ (35)
@@
$\big|\,\hat{\theta}({\hat{p}}_{i,c}\mid\mathbf{x}_{i})-{\hat{p}}_{i,c}\,\big|=\Big|\,\sum_{j\in I_{b}}w_{ij}{\hat{p}}_{j,c}-{\hat{p}}_{i,c}\Big|\,=\big|\,\sum_{j\in I_{b}}w_{ij}({\hat{p}}_{j,c}-{\hat{p}}_{i,c})+\sum_{j\in I_{b}}w_{ij}{\hat{p}}_{i,c}-{\hat{p}}_{i,c}\,\big|=$
$\big|\,\sum_{j\in I_{b}}w_{ij}({\hat{p}}_{j,c}-{\hat{p}}_{i,c})+(\underbrace{\sum_{j\in I_{b}}w_{ij}}_{=1}-1){\hat{p}}_{i,c}\,\big|=\big|\,\sum_{j\in I_{b}}w_{ij}({\hat{p}}_{j,c}-{\hat{p}}_{i,c})\,\big|$
@@
$\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\big|\,\sum_{j\in I_{b}}w_{ij}({\hat{p}}_{j,c}-{\hat{p}}_{i,c})\,\big|\leq\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\underbrace{\sum_{j\in I_{b}}w_{i,j}\big|\,{\hat{p}}_{j,c}-{\hat{p}}_{i,c}\,\big|}_{Z_{i}}$
Again we follow the conditioning assumptions of Appendix A.2. Please note that each coordinate-wise distance satisfies ${\hat{p}}_{j,c}-{\hat{p}}_{i,c}\in[-1,1]$, hence $Z_{i}\in[0,1]$. We can now apply the weighted version of Hoeffding’s inequality to each centered quantity $Z_{i}-\mathbb{E}[Z_{i}]$, conditioning on the kernel weights and features to obtain zero-mean summands. The union bound then yields a simultaneous statement over anchors. We proceed fixing a bin $b$ with index set $I_{b}$ of size $|B_{b}|$. For each anchor $i\in I_{b}$ define the effective sample size associated with the weights:
@@
$\Pr\Bigg(Z_{i}-\mathbb{E}[Z_{i}]\geq\tau\Bigg)\leq\text{exp}\Bigg(-2n_{i}^{\text{eff}}\tau^{2}\Bigg)$
And, by setting $\delta^{\prime}=\delta/nC$ we have:
$\Pr\Bigg(Z_{i}-\mathbb{E}[Z_{i}]\leq\sqrt{\frac{\log\!\bigl(\frac{1}{\delta^{\prime}}\bigr)}{2n_{i}^{\mathrm{eff}}}}\Bigg)\geq 1-\delta^{\prime}$
@@
$Z_{i}-\mathbb{E}[Z_{i}]\leq\sqrt{\frac{\log\!\bigl(\frac{nC}{\delta}\bigr)}{2n_{i}^{\mathrm{eff}}}}$
@@
$\Pr\Bigg(\frac{1}{|B_{b}|}\sum_{i\in I_{b}}Z_{i}\leq\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\mathbb{E}[Z_{i}]+\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\sqrt{\frac{\log\!\bigl(\frac{nC}{\delta}\bigr)}{2n_{i}^{\mathrm{eff}}}}\Bigg)\geq 1-\delta$
We conclude the proof by providing a bound for the expectation in the context of a neural network classifier $\phi(\cdot)$. More precisely, by the Lipschitz continuity of the softmax (Gao and Pavel, 2017),
$\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\sum_{j\in I_{b}}w_{i,j}\,\big|\,{\hat{p}}_{j,c}-{\hat{p}}_{i,c}\,\big|\leq\frac{L}{|B_{b}|}\sum_{i\in I_{b}}\sum_{j\in I_{b}}w_{i,j}\|\phi(\mathbf{x}_{j})-\phi(\mathbf{x}_{i})\|_{1}$